# 2D-OOB: Attributing Data Contribution through Joint Valuation Framework

### v1

@@ Appendix B Additional experimental results
In Section 4.1, we demonstrate that 2D-OOB shows promising performance in identifying cell-level outliers. This section further shows that our result is not sensitive to the selection of hyperparameters. Furthermore, ours generally performs better than 2D-KNN in different settings. We also provide additional results for Section 4.2 and  4.3.
@@ B.1 Additional results for cell-level outlier detection
We have conducted cell-level outlier detection experiments (as in Section 4.1) on three multi-class classification datasets from the UCI Machine Learning repository [19]. As shown in the table, 2D-OOB displays superior detection performance and efficiency.
Table 4: Cell-level outlier experiment results on multi-class classification datasets The average and standard error of the detection AUC and Elapsed Time (in seconds) based on 30 independent experiments are denoted by “average $\pm$ standard error“.
@@
Once we obtain the data valuation scores, an alternative solution approach to determining cell-level attributions involves leveraging feature attribution methods such as SHAP [30]. We explore an additional baseline method building upon this idea: initially, Data-OOB (or any other data valuation method) is computed for the $i$-th data point, denoted as $dv_{i}$. Subsequently, TreeSHAP [31] is fitted, using $dv_{i}$ as the target and the concatenation of $x_{i}$ and $y_{i}$ (denoted as $x_{i}\oplus y_{i}$) as the predictor. The derived local feature attributions are then interpreted as joint valuation results. We refer to this method as ‘two-stage attribution’.
@@ B.2 Additional results for cell fixation experiment
Figure 6: Cell fixation experiment results (test accuracy curves) for 2D-OOB, 2D-KNN and a random baseline. We replace cell values with ground-truth values from the cells with the lowest valuation to the highest valuation. The results from $6$ datasets are displayed. We conduct $30$ independent trials and report the average results. A higher curve indicates better performance. 2D-OOB sets itself apart by its remarkable precision in detecting and rectifying relevant cell outliers.
@@
Following [29], we examine a situation where external information on the ground-truth annotations of outlier cells is not accessible. In this scenario, we address these outliers by substituting them with the average of other cells in the same feature column. This procedure starts by addressing cells with the lowest valuations, based on the hypothesis that correcting these cells is likely to maintain or potentially improve the model’s performance. As depicted in Figure 7, 2D-OOB conforms to this expected trend, demonstrating the effectiveness of our method in joint valuation. Conversely, 2D-KNN fails to show similar performance improvements.
Figure 7: Cell fixation experiment (without ground-truth knowledge) results (test accuracy curves) for 2D-OOB, 2D-KNN and a random baseline. We replace cell values with column mean imputations from the cells with the lowest value to the highest value. The results from $6$ datasets are displayed. We conduct $30$ independent trials and report the average results. A higher curve indicates better performance.
@@ B.3 Additional results for backdoor trigger experiment
We provide additional qualitative examples of backdoor trigger detection experiments in Figure 8.
B.4 Additional results for ablation study
We present the results of ablation study on the number of base learners $B$ and feature subset ratio $K/d$. Specifically, we examine the AUC of the detection curve in the cell-level outlier detection experiment (refer to Table 1).
The number of base learners $B$
When we increase the number of base learners from $500$ to $3000$, the detection AUC for each dataset remains unchanged, as shown in Table 7. This indicates that $1000$ base learners are sufficient to get an equitable joint valuation.
Table 7: Ablation results on the number of base learners $B$. The cell-level outlier detection results (AUC) are examined. Increasing the number of base learners from $1000$ to $3000$ does not yield a notable performance improvement.
Dataset AUC $\uparrow$
$B=500$ $B=1000$ $B=3000$
lawschool 0.86 $\pm$ 0.0035 0.88 $\pm$ 0.0027 0.88 $\pm$ 0.0026
electricity 0.77 $\pm$ 0.0062 0.77 $\pm$ 0.0072 0.77 $\pm$ 0.0070
fried 0.87 $\pm$ 0.0022 0.91 $\pm$ 0.0015 0.91 $\pm$ 0.0014
2dplanes 0.87 $\pm$ 0.0016 0.87 $\pm$ 0.0015 0.87 $\pm$ 0.0015
creditcard 0.72 $\pm$ 0.0025 0.72 $\pm$ 0.0028 0.72 $\pm$ 0.0028
pol 0.78 $\pm$ 0.0022 0.82 $\pm$ 0.0014 0.82 $\pm$ 0.0014
MiniBooNE 0.77 $\pm$ 0.0042 0.77 $\pm$ 0.0058 0.77 $\pm$ 0.0058
jannis 0.78 $\pm$ 0.0045 0.83 $\pm$ 0.0042 0.83 $\pm$ 0.0039
nomao 0.79 $\pm$ 0.0018 0.79 $\pm$ 0.0021 0.79 $\pm$ 0.0020
vehicle_sensIT 0.80 $\pm$ 0.0021 0.81 $\pm$ 0.0014 0.81 $\pm$ 0.0014
gas_drift 0.86 $\pm$ 0.0007 0.86 $\pm$ 0.0010 0.86 $\pm$ 0.0010
musk 0.88 $\pm$ 0.0008 0.88 $\pm$ 0.0008 0.88 $\pm$ 0.0008
Average 0.81 0.83 0.83
Feature subset ratio $K/d$
In addition to $0.50$, We test two additional feature subset ratios $0.25$ and $0.75$. The results in Table 8 suggest that in general, the joint valuation capacity of our method is robust to the choice of feature subset ratio.
Table 8: Ablation results on feature subset ratio $K/d$. The cell-level outlier detection results (AUC) are examined. Our method’s joint valuation capacity remains relatively stable regardless of the selected feature subset ratio.
Dataset AUC $\uparrow$
$K/d=0.25$ $K/d=0.50$ $K/d=0.75$
lawschool 0.86$\pm$ 0.0026 0.88$\pm$ 0.0027 0.88$\pm$ 0.0024
electricity 0.79$\pm$ 0.0070 0.77$\pm$ 0.0072 0.73$\pm$ 0.0070
fried 0.86$\pm$ 0.0024 0.91$\pm$ 0.0015 0.89$\pm$ 0.0007
2dplanes 0.82$\pm$ 0.0015 0.87$\pm$ 0.0015 0.88$\pm$ 0.0014
creditcard 0.73$\pm$ 0.0029 0.72$\pm$ 0.0028 0.71$\pm$ 0.0028
pol 0.66$\pm$ 0.0031 0.82$\pm$ 0.0014 0.82$\pm$ 0.0014
MiniBooNE 0.78$\pm$ 0.0076 0.77$\pm$ 0.0058 0.77$\pm$ 0.0049
jannis 0.84$\pm$ 0.0035 0.83$\pm$ 0.0042 0.82$\pm$ 0.0043
nomao 0.79$\pm$ 0.0019 0.79$\pm$ 0.0021 0.78$\pm$ 0.0021
vehicle_sensIT 0.81$\pm$ 0.0014 0.81$\pm$ 0.0014 0.80$\pm$ 0.0015
gas_drift 0.88$\pm$ 0.0009 0.86$\pm$ 0.0010 0.86$\pm$ 0.0009
musk 0.89$\pm$ 0.0008 0.88$\pm$ 0.0008 0.88$\pm$ 0.0008
Average 0.81 0.83 0.82
Figure 8: Qualitative results on more datasets for the backdoor trigger detection experiment. The first two images originate from the class “airplane” while relabeled as “automobile”. The latter two images originate from the class “frog” while relabeled as “horse”.
@@ B.4 Additional results for ablation study
For simplicity, we denote $\phi^{\mathrm{OOB}}_{i}(S)$ as $\phi_{i}(S)$ and $\psi^{\mathrm{2D-OOB}}_{ij}$ as $\psi_{ij}$ in the proof. With the set of subsets $\mathcal{S}:=\{S\subseteq[d]\}$ and the definition Data-OOB $\phi_{i}(S)$ for all $i\in[n]$, where $S$ is a feature subset, we denote the cardinality of $\mathcal{S}$ as $L:=|\mathcal{S}|=2^{d}$. Let $\gamma_{b}$ be a weight vector $\gamma_{b}:=(\gamma_{b1},\ldots,\gamma_{bL})$ for all $b\in[B]$, where $\gamma_{bl}\in\{0,1\}$ and $\gamma_{bl}=1$ indicates the $l$-th subset is used in the $b$-th weak learner. With $\{w_{b},\gamma_{b},\hat{f}_{b}\}_{b=1}^{B}$, we can also denote the $i$-th Data-OOB on $l$-th feature subset $S_{l}$ as
@@
where $\alpha_{i,j,l}\propto\mathds{1}(j\in S_{l})\sum_{b=1}^{B}\mathds{1}(w_{bi}=0)\mathds{1}(\gamma_{bl}=1),\forall i\in[n],j\in[d],l\in[L]$ and $\sum_{l=1}^{L}\alpha_{i,j,l}=1$. Define $P_{i}(S_{l}|j\in S_{l},\{w_{bi}\}_{b=1}^{B})=\alpha_{i,j,l}$, which specifies an empirical distribution of the feature subset $S$, conditioned on $j\in S$, in relation to the bootstrap sampling process. Here, $\mathds{1}(j\in S_{l})$ implies the distribution is conditioned on the presence of the $j$-th feature within the feature subset $S_{l}$. $w_{bi}$ indicates whether the $i$-th sample is out-of-bag in the $b$-th bootstrap, and $\gamma_{bl}$ indicates whether the $l$-th feature subset is selected in the $b$-th weak learner. Thus, the point mass is determined by the sampling process.
@@ Appendix D Data valuation experiment
$\psi^{data}_{i}:=\frac{1}{d}\sum_{j=1}^{d}\psi^{\mathrm{2D-OOB}}_{ij},$ (5)
Proof.
Based on definition of 2D-OOB-Data, for $i\in[n]$,
where $\alpha_{i,j,l}$ is defined in Appendix C. We have $\sum_{l=1}^{L}(\frac{1}{d}\sum_{j=1}^{d}\alpha_{i,j,l})=\frac{1}{d}\sum_{j=1}^{d}\sum_{l=1}^{L}\alpha_{i,j,l}=1$. Denote $P_{i}(S_{l}|\{w_{bi}\}_{b=1}^{B})=\frac{1}{d}\sum_{j=1}^{d}\alpha_{i,j,l}$, which induces the empirical expectation of Data-OOB with respect to $S_{l}$. ∎
@@
Proposition D.1 indicates 2D-OOB-data $\psi^{data}_{i}$ can be expressed as the average Data-OOB value for the $i$-th data point. As a result, 2D-OOB-data is expected to inherit the advanced ability of Data-OOB in terms of data valuation, as will be empirically examined in the Appendix D.
@@
Following the standard protocol in Kwon and Zou [22, 23] and Jiang et al. [16], we randomly select $10$% of the data points and change its label to the other class. For joint valuation methods, we calculate the valuation of each cell and perform the marginalization over features to obtain the data valuation scores. For the baseline methods, we further incorporate several state-of-the-art data valuation methods including DataShapley [12], KNNShapley [15], DataBanzhaf [46], LAVA [18], and Data-OOB [23]. Implementation details are listed below. To guarantee a fair comparison, we also employ the decision tree as the base model in DataShapley and DataBanzhaf. Mislabeled data detection and data removal experiment are examined based on this setting. We adopt the same $12$ datasets as outlined in Section 4.1.
@@
We use a Monte Carlo-based algorithm. The Gelman-Rubin statistics is computed to determine the termination criteria of the algorithm. Following Jiang et al. [16], We adopt the threshold to be $1.05$. To ensure a fair comparison with the proposed method, we employ the decision tree model for the utility evaluation.
@@
We adopt the implementation from Jiang et al. [16]. We employ the decision tree model and set “the number of models to train” to $1000$.
@@ D.1 Mislabeled data detection
We calculate the precision-recall curve by comparing the actual annotations, which denote whether data points are mislabeled, against the data valuation scores computed by different methods. Mislabeled data typically have a detrimental impact on model performance. Therefore, data points that receive a lower valuation score are regarded as having a higher chance of being mislabeled. We then determine AUCPR (the AUC of the precision-recall curve) as a quantitative metric to assess the detection efficacy.

### v2

@@ B.1 Additional results for cell-level outlier detection
We conducted cell-level outlier detection experiments (as described in Section 4.1) on three multi-class classification datasets from the UCI Machine Learning repository [19]. As shown in the Table 5, 2D-OOB displays superior detection performance and efficiency.
Table 5: Cell-level outlier detection results on multi-class classification datasets. The average and standard error of the AUC and run-time (in seconds) based on $30$ independent experiments are denoted by “average $\pm$ standard error”.
@@
Once we obtain the data valuation scores, an alternative solution approach to determining cell-level attributions involves leveraging feature attribution methods such as SHAP [31]. We explore an additional baseline method building upon this idea: initially, Data-OOB (or any other data valuation method) is computed for the $i$-th data point, denoted as $dv_{i}$. Subsequently, TreeSHAP [32] is fitted, using $dv_{i}$ as the target and the concatenation of $x_{i}$ and $y_{i}$ (denoted as $x_{i}\oplus y_{i}$) as the predictor. The derived local feature attributions are then interpreted as joint valuation results. We refer to this method as “two-stage attribution”.
@@ B.2 Additional results for cell fixation experiment
Figure 6: Cell fixation experiment results (test accuracy curves) for 2D-OOB, 2D-KNN and a random baseline. We replace cell values with ground-truth values from the cells with the lowest valuation to the highest valuation. The results from $6$ additional datasets are displayed. We conduct $30$ independent trials and report the average results. A higher curve indicates better performance. 2D-OOB sets itself apart by its remarkable precision in detecting and rectifying relevant cell outliers.
@@
Following Liu et al. [30], we examine a situation where external information on the ground-truth annotations of outlier cells is not accessible. In this scenario, we address these outliers by substituting them with the average of other cells in the same feature column. This procedure starts by addressing cells with the lowest valuations, based on the hypothesis that correcting these cells is likely to maintain or potentially improve the model’s performance. As depicted in Figure 7, 2D-OOB conforms to this expected trend, demonstrating the effectiveness of our method in joint valuation. Conversely, 2D-KNN fails to show similar performance improvements.
Figure 7: Cell fixation experiment (without ground-truth knowledge) results (test accuracy curves) for 2D-OOB, 2D-KNN and a random baseline. We replace cell values with column mean imputations from cells with the lowest value to the highest value. We report the average results of $30$ independent trials from $12$ datasets. A higher curve indicates better performance.
@@ B.3 Additional results for backdoor trigger experiment
We provide additional qualitative examples of the backdoor trigger detection experiment in Figure 8.
Figure 8: Qualitative results on more datasets for the backdoor trigger detection experiment. The first two images are from the class “airplane” but have been relabeled as “automobile”, while the latter two images are from the class “frog” and have been relabeled as “horse”.
@@
For simplicity, we denote $\phi^{\mathrm{OOB}}_{i}(S)$ as $\phi_{i}(S)$ and $\psi^{\mathrm{2D-OOB}}_{ij}$ as $\psi_{ij}$ in the proof. Let $\mathcal{S}:=\{S\subseteq[d]\}$ represent the set of all feature subsets, where $S$ is a feature subset. We denote the cardinality of $\mathcal{S}$ as $L:=|\mathcal{S}|=2^{d}$. Let $\boldsymbol{\gamma}_{b}$ be a weight vector $\boldsymbol{\gamma}_{b}:=(\gamma_{b1},\ldots,\gamma_{bL})$ for all $b\in[B]$, where $\gamma_{bl}\in\{0,1\}$ and $\gamma_{bl}=1$ indicates the $l$-th subset is used in the $b$-th weak learner. With $\{\mathbf{w}_{b},\boldsymbol{\gamma}_{b},\hat{f}_{b}\}_{b=1}^{B}$, we can denote the $i$-th Data-OOB on the $l$-th feature subset $S_{l}$ as
@@
where $\alpha_{i,j,l}\propto\mathds{1}(j\in S_{l})\sum_{b=1}^{B}\mathds{1}(w_{bi}=0)\mathds{1}(\gamma_{bl}=1),\forall i\in[n],j\in[d],l\in[L]$ and $\sum_{l=1}^{L}\alpha_{i,j,l}=1$. Define $P_{i}(S_{l}|j\in S_{l},\{w_{bi}\}_{b=1}^{B})=\alpha_{i,j,l}$, which specifies an empirical distribution of the feature subset $S$, conditioned on $j\in S$ and the bootstrap sampling process. Here, $\mathds{1}(j\in S_{l})$ indicates that the distribution is conditioned on the inclusion of the $j$-th feature within the feature subset $S_{l}$. $w_{bi}$ indicates whether the $i$-th sample is out-of-bag in the $b$-th bootstrap, and $\gamma_{bl}$ indicates whether the $l$-th feature subset is selected in the $b$-th weak learner. Thus, the point mass is determined by the sampling process, leading to:
$\psi_{ij}=\mathbb{E}_{\hat{F}_{S}}[\phi_{i}(S)\mid j\in S].$
@@ Appendix D Data valuation experiment
$\psi^{data}_{i}:=\frac{1}{d}\sum_{j=1}^{d}\psi^{\mathrm{2D-OOB}}_{ij}.$ (5)
@@
Proof.
Based on the definition of 2D-OOB-Data, for $i\in[n]$,
where $\alpha_{i,j,l}$ is defined in Appendix C. We have $\sum_{l=1}^{L}(\frac{1}{d}\sum_{j=1}^{d}\alpha_{i,j,l})=\frac{1}{d}\sum_{j=1}^{d}\sum_{l=1}^{L}\alpha_{i,j,l}=1$. Denote $\mathbb{P}_{i}(S_{l}|\{w_{bi}\}_{b=1}^{B})=\frac{1}{d}\sum_{j=1}^{d}\alpha_{i,j,l}$, which induces the empirical expectation of Data-OOB with respect to $S_{l}$. ∎
Proposition D.1 indicates 2D-OOB-data $\psi^{data}_{i}$ can be expressed as the average Data-OOB value for the $i$-th data point. As a result, 2D-OOB-data is expected to inherit the advanced ability of Data-OOB in terms of data valuation, as will be empirically examined next.
@@
Following the standard protocol in Kwon and Zou [23, 24] and Jiang et al. [16], we randomly select $10$% of the data points and flip its label to the other class. For joint valuation methods, we calculate the valuation of each cell and perform the marginalization over features to obtain the data valuation scores. Mislabeled data detection and data removal experiments are examined based on this setting. For the baseline methods, we further incorporate several state-of-the-art data valuation methods including DataShapley [12], KNNShapley [15], DataBanzhaf [48], LAVA [18], and Data-OOB [24]. Implementation details are listed below. To guarantee a fair comparison, we also employ the decision tree as the base model in DataShapley and DataBanzhaf. We adopt the same $12$ datasets as outlined in Section 4.1.
@@
We use a Monte Carlo-based algorithm. The Gelman-Rubin statistics is computed to determine the termination criteria of the algorithm. Following Jiang et al. [16], We adopt the threshold to be $1.05$.
@@
We adopt the implementation from Jiang et al. [16]. We set “the number of models to train” as $1000$.
@@ D.1 Mislabeled data detection
We calculate the precision-recall curve by comparing the actual annotations, which denote whether data points are mislabeled, against the data valuation scores computed by different methods. Mislabeled data typically have a detrimental impact on model performance. Therefore, data points that receive a lower valuation score are considered to have a higher chance of being mislabeled. We then determine AUCPR (the AUC of the precision-recall curve) as a quantitative metric to assess the detection efficacy.